OpenStudy (anonymous):

please explain @ganeshie8

OpenStudy (akashdeepdeb):



OpenStudy (akashdeepdeb):

Now, here the limit of x approaches 4 from the left side. So basically, the value should be = -1. \[\lim_{x \rightarrow 4^-} \frac{|x-4|}{x-4} = -1\]

OpenStudy (anonymous):

thanks!

OpenStudy (phi):

to add to Akash the \( x \rightarrow 4^-\) means x is less than 4 and approaches (but *never* reaches 4) example 3.9 (to make it concrete) | 3.9 - 4 | / (3.9- 4) = |-0.1|/ -0.1 = 0.1/-0.1 = -1 As x gets closer to 4 we can replace 0.1 with a *tiny* number \( \epsilon\) \[ \frac{\epsilon}{- \epsilon} =-1 \] the important idea is \( \epsilon \) is *never* allowed to be zero, so we can always divide to get -1